# neurips_2025.bib Edge-ASR: Towards Low-Bit Quantization of Automatic Speech Recognition Models

### v1

@@ 3.3 Overall Comparison
Table 8 provides an overall comparisons of quantized edge-ASR models across key deployment-relevant metrics. Model size refers primarily to the total memory required to store model parameters. This has a significant impact on both latency and power consumption, particularly when weights cannot fit on-chip, resulting in frequent external memory accesses. Memory I/O captures the data movement involved in loading inputs and weights, and storing outputs for each computational operation. Since memory I/O can dominate both latency and energy usage, minimizing memory access is critical for achieving energy-efficient and low-latency inference. Bit Operations (BOPs) measures the total computational cost, taking into account both the number of MAC (multiply-accumulate) operations and data bit-width. Higher BOPs correspond to increased computing energy. Therefore, reducing bit-width not only lowers model size, but also significantly decreases the energy consumed per operation. For ASR models, the encoder is constrained by activation size, where memory I/O dominates the energy usage, while the decoder is constrained by weight size. Therefore, for low-power use cases, configurations such as w8-a8 for encoder, and w4-a16 for decoder would be considered for edge application. Based on the accuracy criteria and hardware specifications, users can choose proper models and bit-width configurations upon deployment.
@@ 4 Limitations
While our benchmark provides a comprehensive assessment of post-training quantization for edge-ASR models, several limitations remain. Model diversity. Our study focuses on two transformer-based ASR model families, Whisper and Moonshine, due to their compact model sizes, leading performance, and suitability for edge deployment. Future work will expand the edge-ASR models to include additional architectures (e.g., RNN-based, state space models (SSMs), Conformer variants and Canary), to broaden the applicability of our findings. Algorithm coverage. We evaluate eight state-of-the-art PTQ methods covering transformation, reconstruction and rounding-based strategies. Emerging techniques, such as rotation-based quantization, are not yet supported in our current toolkit. Incorporating these newer algorithms will enrich the benchmark and offer more effective options for aggressive model compression. Framework extensibility. Although we extend the LLM compression toolkit to support ASR models and a diverse set of quantization algorithms, many other factors remain unexplored (e.g., dynamic quantization, hardware-specific quantization strategies). Future enhancements will expand our framework to include these dimensions, enabling comprehensive support for end-to-end model optimization, conversion and deployment on a wide range of edge platforms. Addressing these limitations in future work will help guide the development of more efficient and accurate quantized ASR systems tailored for resource-constrained environments.
@@ 5 Conclusion
We present the first systematic benchmark of eight post-training quantization (PTQ) methods on two Transformer-based edge-ASR model families (Whisper and Moonshine), evaluated across seven diverse datasets. Through comprehensive analysis of critical quantization factors, including bit-width selection, ultra-low bit quantization, granularity, weight clipping, and symmetric versus asymmetric quantization, along with ablation studies on calibration data, we characterize the trade-offs between model accuracy, model size, memory footprint, I/O overhead, and computational cost in the fiexed-point domain, establishing a quantized standard for the open ASR leaderboard. Our findings show that advanced PTQ strategies enable viable quantization down to -bit weights. By building an extension to the LLM compression toolkit to support ASR architectures, we provide a workflow that enables rapid, reproducible quantization of edge-ASR models. These insights and tools serve to bridge the gap between model training and edge deployment. Future work will broaden the model zoo to include a wider range of ASR architectures, incorporate emerging quantization techniques, and integrate deployment pipelines for edge platforms, faciliating efficient and accurate quantized ASR systems on low-power, always-on edge devices.

### v2

@@ 3.3 Overall Comparison
Table 8 showcases the overall comparisons of quantized edge-ASR models across key deployment-relevant metrics. Model size refers primarily to the total memory required to store model parameters. This has a significant impact on both latency and power consumption, particularly when weights cannot fit on-chip, resulting in frequent external memory accesses. Memory I/O captures the data movement involved in loading inputs and weights, and storing outputs for each computational operation. Since memory I/O can dominate both latency and energy usage, minimizing memory access is critical for achieving energy-efficient and low-latency inference. Bit Operations (BOPs) measures the total computational cost, taking into account both the number of MAC (multiply-accumulate) operations and data bit-width. Higher BOPs correspond to increased computing energy. Therefore, reducing bit-width not only lowers model size, but also significantly decreases the energy consumed per operation. For ASR models, the encoder is constrained by activation size, where memory I/O dominates the energy usage, while the decoder is constrained by weight size. Therefore, for low-power use cases, configurations such as w8-a8 for encoder, and w4-a16 for decoder would be considered for edge application. Based on the accuracy criteria and hardware specifications, users can choose appropriate models and bit-width configurations upon deployment.
@@ 4 Limitations
While our benchmark provides a comprehensive assessment of post-training quantization for edge-ASR models, several limitations remain. Model diversity. Our study focuses on two transformer-based ASR model families, Whisper and Moonshine, due to their compact model sizes, leading performance, and suitability for edge deployment. Future work will expand the edge-ASR models to include additional architectures (e.g., RNN-based, state space models (SSMs), Conformer variants and Canary), to broaden the applicability of our findings. Algorithm coverage. We evaluate eight state-of-the-art PTQ methods covering transformation, reconstruction and rounding-based strategies. Emerging techniques, such as rotation-based quantization, are not yet supported in our current toolkit. Incorporating these newer algorithms will enrich the benchmark and offer more effective options for aggressive model compression. Framework extensibility. Although we extend the LLM compression toolkit to support ASR models and a diverse set of quantization algorithms, many other factors remain unexplored (e.g., dynamic quantization, hardware-specific quantization strategies). Future enhancements will expand our framework to include these dimensions, enabling comprehensive support for end-to-end model optimization, conversion and deployment on a wide array of edge platforms. Addressing these limitations in future work will help guide the development of more efficient and accurate quantized ASR systems tailored for resource-constrained environments.
@@ 5 Conclusion
We present the first systematic benchmark of eight post-training quantization (PTQ) methods on two Transformer-based edge-ASR model families (Whisper and Moonshine), evaluated across seven diverse datasets. Through comprehensive analysis of critical quantization factors, including bit-width selection, ultra-low bit quantization, granularity, weight clipping, and symmetric versus asymmetric quantization, along with ablation studies on calibration data, we characterize the trade-offs between model accuracy, model size, memory footprint, I/O overhead, and computational cost in the fixed-point domain, establishing a quantized standard for the open ASR leader-board. Our findings show that advanced PTQ strategies enable viable quantization down to -bit weights. By building an extension to the LLM compression toolkit to support ASR architectures, we provide a workflow that enables rapid, reproducible quantization of edge-ASR models. These insights and tools serve to bridge the gap between model training and edge deployment. Future work will broaden the model zoo to include a wide range of ASR architectures, incorporate emerging quantization techniques, and integrate deployment pipelines for edge platforms, facilitating efficient and accurate quantized ASR systems on low-power, always-on edge devices.